OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? help! x2 + 17x

OpenStudy (jim766):

17/2 = 8.5 now square it 8.5^2 = 72.25

OpenStudy (anonymous):

5220.06

OpenStudy (anonymous):

0625

OpenStudy (anonymous):

but it says a number

OpenStudy (jim766):

the number you add to the polynomial is 72.25

OpenStudy (jim766):

x^2+17x+72.25 = (x+8.5)^2

OpenStudy (anonymous):

so is 72.25

OpenStudy (jim766):

right

OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? x2 - 10x

OpenStudy (anonymous):

this is 25 right

OpenStudy (jim766):

right take 10, divide by 2 then square

OpenStudy (anonymous):

is 25= 5

OpenStudy (anonymous):

so is 25 right

OpenStudy (anonymous):

or 5

OpenStudy (jim766):

yes

OpenStudy (jim766):

25

OpenStudy (jim766):

10/2 = 5, 5^2 = 25

OpenStudy (anonymous):

oh god thanks i got this thanks for the help!

OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? x2 - 28x whbu this one

OpenStudy (anonymous):

x^2-28

OpenStudy (anonymous):

you cant divide

OpenStudy (jim766):

-28/2 = -14, -14^2 = 196

OpenStudy (anonymous):

196

OpenStudy (jim766):

right

OpenStudy (anonymous):

even though the x^2 is square u could still divide it right

OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? x2 - 19x

OpenStudy (jim766):

do you mean...if the number is negative can you still divide it? yes ,

OpenStudy (jim766):

-19/2 =

OpenStudy (anonymous):

-9.5

OpenStudy (jim766):

now what is 9.5^2

OpenStudy (anonymous):

90.25

OpenStudy (jim766):

right, the negative sign does not matter...because when you square something, the answer is always positive

OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? x2 + 7x

OpenStudy (jim766):

7/2 =

OpenStudy (anonymous):

3.5

OpenStudy (jim766):

now square it

OpenStudy (anonymous):

12.25

OpenStudy (jim766):

right

OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? this one 121 x2 - 22x

OpenStudy (jim766):

perfect

OpenStudy (anonymous):

i got it thanks

OpenStudy (jim766):

yw

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

i got a 92

OpenStudy (anonymous):

on the quiz

OpenStudy (jim766):

nice!